Suppose that the annual interest rate is 2.5 percent in Korea and 4.2 percent in Germany, and tha...

  1. Suppose that the annual interest rate is 2.5 percent in Korea and 4.2 percent in Germany, and that the spot exchange rate is Won1933.2/€ and the forward exchange rate, with one-year maturity, is W1915.5/€. Assume that a trader can borrow up to €2,000,000 or Won3,866,400,000.

  1. Does the interest rate parity hold? Show your work.
  2. Is there an arbitrage opportunity? (covered interest arbitrage)
  3. If there is an arbitrage opportunity, what steps should we take in order to make an arbitrage profit?
  4. What will be our maximum profit?
Answer #1

As per interest rate parity, forward rate = spot rate*(1+Interest rateKorea)/(1+Interest rate Germany)

= 1933.2*(1+0.025)/(1+0.042)

= Won1,901.66/Euro

Since the actual forward rate is different from fair forward rate, interest rate parity DOES NOT HOLD

b.Yes, since interest rate parity does not hold

c.Borrow Won3,866,400,000

Convert into Euro at Spot Rate = 3,866,400,000/1933.2

= Euro 2,000,000

Invest and get 2,000,000(1+0.042) = Euro 2,084,000

Convert back into Won at forward rate = 2,084,000*1,915.5 = Won3,991,902,000

Pay back loan 3,866,400,000*(1.025) = Won 3,963,060,000

Arbitrage Profit = Won 28,842,000

D. profit = Won 28,842,000